OpenStudy (anonymous):

how can I do this? 2/x + 5/25 = 3/5

OpenStudy (anonymous):

Do I subtract the 3/25 over and flip both sides?

OpenStudy (mathstudent55):

The best way, at least for me, is multiply both sides by the least common denominator of the fractions and get rid of denominators. You have denominators x, 5, 25. What is the LCD?

OpenStudy (anonymous):

would 5 be?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

The 3 denominators are 5, x, 5(5). The LCD is 5(5)(x) or 25x

OpenStudy (mathstudent55):

Now multiply both sides by 25x, and all denominators will cancel out.

OpenStudy (anonymous):

so from 2/x + 5/25 = 3/5 I would get 10/25+ 5/25= 15/25?

OpenStudy (anonymous):

so x=15?

OpenStudy (mathstudent55):